Given the following demand function and total cost function as:

Q = 250-5P and TC = 0.5Q2 + 5Q + 125 

Calculate the equilibrium price, Equilibrium Quantity and the total profit

A. under perfectly competitive market structure

B. Under monopolistic competitive market


1
Expert's answer
2022-01-02T18:19:25-0500

A. Under perfectly competitive market structure in equilibrium D = MC, so:

"MC = TC'(Q) = Q + 5,"

50 - 0.2Q = Q + 5,

Q = 37.5 units,

"P = 50 - 0.2*37.5 = 42.5."


B. Under monopolistic competitive market in equilibrium MR = MC, so:

"MR = TR'(Q) = 50 - 0.4Q,"

50 - 0.4Q = Q + 5,

Q = 32.14 units,

"P = 50 - 0.2*32.14 = 43.57."